7. Determine the lateral area of a cylinder of height 300mm and base radius of 100mm.
1
Expert's answer
2020-12-10T20:26:16-0500

Consider a right circular cylinder with radius of base "r" and height "h"



The lateral area of the cylinder is "F=2\pi rh"

Given "h=300\ mm=30\ cm, r=100\ mm=10\ cm"

Then


"F=2\pi(10\ cm)(30\ cm)=600\pi\ cm^2\approx1885\ cm^2"